\begin{thebibliography}{10}

\bibitem{idvBook_v2}
M.~O. Ward, G.~Grinstein, and D.~Keim, {\em Interactive Data Visualization:
  Foundations, Techniques, and Applications, Second Edition}.
\newblock A. K. Peters, Ltd., 2015.

\bibitem{bs16}
N.~Bikakis and T.~Sellis, ``{E}xploration and {V}isualization in the {W}eb of
  {B}ig {L}inked {D}ata: {A} {S}urvey of the {S}tate of the {A}rt,'' in {\em
  6th Intl. Workshop on Linked Web Data Management (LWDM)}, 2016.

\bibitem{Shneiderman08}
B.~Shneiderman, ``{E}xtreme {V}isualization: {S}queezing a {B}illion {R}ecords
  into a {M}illion {P}ixels,'' in {\em {ACM} Conference on Management of Data
  ({SIGMOD})}, 2008.

\bibitem{HeerK12b}
J.~Heer and S.~Kandel, ``{I}nteractive {A}nalysis of {B}ig {D}ata,'' {\em {ACM}
  Crossroads}, vol.~19, no.~1, 2012.

\bibitem{MortonBGM14}
K.~Morton, M.~Balazinska, D.~Grossman, and J.~D. Mackinlay, ``{S}upport the
  {D}ata {E}nthusiast: {C}hallenges for {N}ext-generation {D}ata-analysis
  {S}ystems,'' {\em VLDB Endowment (PVLDB)}, vol.~7, no.~6, 2014.

\bibitem{WuBM14}
E.~Wu, L.~Battle, and S.~R. Madden, ``{T}he {C}ase for {D}ata {V}isualization
  {M}anagement {S}ystems,'' {\em VLDB Endowment (PVLDB)}, vol.~7, no.~10, 2014.

\bibitem{GGL15}
P.~Godfrey, J.~Gryz, and P.~Lasek, ``{I}nteractive {V}isualization of {L}arge
  {D}ata {S}ets,'' {\em {IEEE} Transactions on Knowledge and Data Engineering
  ({TKDE})}, vol.~28, no.~8, 2016.

\bibitem{IdreosPC15}
S.~Idreos, O.~Papaemmanouil, and S.~Chaudhuri, ``{O}verview of {D}ata
  {E}xploration {T}echniques,'' in {\em {ACM} Conference on Management of Data
  ({SIGMOD})}, 2015.

\bibitem{HeerS12}
J.~Heer and B.~Shneiderman, ``{I}nteractive {D}ynamics for {V}isual
  {A}nalysis,'' {\em Commun. {ACM}}, vol.~55, no.~4, 2012.

\bibitem{FisherPDs12}
D.~Fisher, I.~O. Popov, S.~M. Drucker, and M.~C. Schraefel, ``{T}rust {M}e,
  {I}'m {P}artially {R}ight: {I}ncremental {V}isualization {L}ets {A}nalysts
  {E}xplore {L}arge {D}atasets {F}aster,'' in {\em Conference on Human Factors
  in Computing Systems (CHI)}, 2012.

\bibitem{ParkCM15}
Y.~Park, M.~J. Cafarella, and B.~Mozafari, ``{V}isualization-aware {S}ampling
  for {V}ery {L}arge {D}atabases,'' in {\em {IEEE} Intl. Conf. on Data
  Engineering ({ICDE})}, 2016.

\bibitem{AgarwalMPMMS13}
S.~Agarwal, B.~Mozafari, A.~Panda, H.~Milner, S.~Madden, and I.~Stoica,
  ``{B}linkdb: {Q}ueries with {B}ounded {E}rrors and {B}ounded {R}esponse
  {T}imes on {V}ery {L}arge {D}ata,'' in {\em European Conference on Computer
  Systems (EuroSys)}, 2013.

\bibitem{ImVM13}
J.~Im, F.~G. Villegas, and M.~J. McGuffin, ``{V}isreduce: {F}ast and
  {R}esponsive {I}ncremental {I}nformation {V}isualization of {L}arge
  {D}atasets,'' in {\em {IEEE} Conf. on Big Data ({BigData})}, 2013.

\bibitem{BattleSC13}
L.~Battle, M.~Stonebraker, and R.~Chang, ``{D}ynamic {R}eduction of {Q}uery
  {R}esult {S}ets for {I}nteractive {V}isualizaton,'' in {\em {IEEE} Conf. on
  Big Data ({BigData})}, 2013.

\bibitem{EF10}
N.~Elmqvist and J.~Fekete, ``{H}ierarchical {A}ggregation for {I}nformation
  {V}isualization: {O}verview, {T}echniques, and {D}esign {G}uidelines,'' {\em
  {IEEE} Trans. Vis. Comput. Graph. ({TVCG})}, vol.~16, no.~3, 2010.

\bibitem{bsps15}
N.~Bikakis, G.~Papastefanatos, M.~Skourla, and T.~Sellis, ``{A} {H}ierarchical
  {A}ggregation {F}ramework for {E}fficient {M}ultilevel {V}isual {E}xploration
  and {A}nalysis,'' {\em Semantic Web Journal}, vol.~8, no.~1, 2017.

\bibitem{JugelJM15}
U.~Jugel, Z.~Jerzak, G.~Hackenbroich, and V.~Markl, ``{VDD}a: {A}utomatic
  {V}isualization-driven {D}ata {A}ggregation in {R}elational {D}atabases,''
  {\em Journal on Very Large Data Bases ({VLDBJ})}, 2015.

\bibitem{LiuJH13}
Z.~Liu, B.~Jiang, and J.~Heer, ``\emph{imMens}: {R}eal-time {V}isual {Q}uerying
  of {B}ig {D}ata,'' {\em Comput. Graph. Forum ({CGF})}, vol.~32, no.~3,
  pp.~421--430, 2013.

\bibitem{LinsKS13}
L.~D. Lins, J.~T. Klosowski, and C.~E. Scheidegger, ``{N}anocubes for
  {R}eal-time {E}xploration of {S}patiotemporal {D}atasets,'' {\em {IEEE}
  Trans. Vis. Comput. Graph. ({TVCG})}, vol.~19, no.~12, 2013.

\bibitem{GodfreyGLR16}
P.~Godfrey, J.~Gryz, P.~Lasek, and N.~Razavi, ``{V}isualization through
  {I}nductive {A}ggregation,'' in {\em Intl. Conf. on Extending Database
  Technology ({EDBT})}, 2016.

\bibitem{Shneiderman96}
B.~Shneiderman, ``{T}he {E}yes {H}ave {I}t: {A} {T}ask by {D}ata {T}ype
  {T}axonomy for {I}nformation {V}isualizations,'' in {\em IEEE Symposium on
  Visual Languages}, 1996.

\bibitem{RodriguesTPTTF13}
J.~F.~R. Jr., H.~Tong, J.~Pan, A.~J.~M. Traina, C.~T. Jr., and C.~Faloutsos,
  ``{L}arge {G}raph {A}nalysis in the {GM}ine {S}ystem,'' {\em {IEEE}
  Transactions on Knowledge and Data Engineering ({TKDE})}, vol.~25, no.~1,
  2013.

\bibitem{AbelloHK06}
J.~Abello, F.~van Ham, and N.~Krishnan, ``{ASK}-graphview: {A} {L}arge {S}cale
  {G}raph {V}isualization {S}ystem,'' {\em {IEEE} Trans. Vis. Comput. Graph.
  ({TVCG})}, vol.~12, no.~5, 2006.

\bibitem{Auber04}
D.~Auber, ``{T}ulip - a {H}uge {G}raph {V}isualization {F}ramework,'' in {\em
  Graph Drawing Software}, 2004.

\bibitem{TominskiAS09}
C.~Tominski, J.~Abello, and H.~Schumann, ``{C}gv - an {I}nteractive {G}raph
  {V}isualization {S}ystem,'' {\em Computers {\&} Graphics}, vol.~33, no.~6,
  2009.

\bibitem{LinCTWKC13}
Z.~Lin, N.~Cao, H.~Tong, F.~Wang, U.~Kang, and D.~H.~P. Chau, ``{D}emonstrating
  {I}nteractive {M}ulti-resolution {L}arge {G}raph {E}xploration,'' in {\em
  ICDM Workshops}, 2013.

\bibitem{ZinsmaierBDS12}
M.~Zinsmaier, U.~Brandes, O.~Deussen, and H.~Strobelt, ``{I}nteractive
  {L}evel-of-detail {R}endering of {L}arge {G}raphs,'' {\em {IEEE} Trans. Vis.
  Comput. Graph. ({TVCG})}, vol.~18, no.~12, 2012.

\bibitem{ArchambaultMA08}
D.~Archambault, T.~Munzner, and D.~Auber, ``{G}rouseflocks: {S}teerable
  {E}xploration of {G}raph {H}ierarchy {S}pace,'' {\em {IEEE} Trans. Vis.
  Comput. Graph. ({TVCG})}, vol.~14, no.~4, 2008.

\bibitem{SundaraAKDWCS10}
S.~Sundara, M.~Atre, V.~Kolovski, S.~Das, Z.~Wu, E.~I. Chong, and
  J.~Srinivasan, ``{V}isualizing {L}arge-scale {R}df {D}ata {U}sing {S}ubsets,
  {S}ummaries, and {S}ampling in {O}racle,'' in {\em {IEEE} Intl. Conf. on Data
  Engineering ({ICDE})}, 2010.

\bibitem{BikakisLKG16}
N.~Bikakis, J.~Liagouris, M.~Krommyda, G.~Papastefanatos, and T.~Sellis,
  ``{G}raphvizdb: {A} {S}calable {P}latform for {I}nteractive {L}arge {G}raph
  {V}isualization,'' in {\em {IEEE} Intl. Conf. on Data Engineering ({ICDE})},
  2016.

\bibitem{Gansner2011}
E.~R. Gansner, Y.~Hu, S.~C. North, and C.~E. Scheidegger, ``{M}ultilevel
  {A}gglomerative {E}dge {B}undling for {V}isualizing {L}arge {G}raphs,'' in
  {\em {IEEE} Pacific Visualization Symposium (PacificVis)}, 2011.

\bibitem{Ersoy2011}
O.~Ersoy, C.~Hurter, F.~V. Paulovich, G.~Cantareiro, and A.~Telea,
  ``{S}keleton-based {E}dge {B}undling for {G}raph {V}isualization,'' {\em
  {IEEE} Trans. Vis. Comput. Graph. ({TVCG})}, vol.~17, no.~12, 2011.

\bibitem{Phan2005}
D.~Phan, L.~Xiao, R.~B. Yeh, P.~Hanrahan, and T.~Winograd, ``{F}low {M}ap
  {L}ayout,'' in {\em {IEEE} Symposium on Information Visualization (InfoVis)},
  2005.

\bibitem{Lambert2010}
A.~Lambert, R.~Bourqui, and D.~Auber, ``{W}inding {R}oads: {R}outing {E}dges
  into {B}undles,'' {\em Comput. Graph. Forum ({CGF})}, vol.~29, no.~3, 2010.

\bibitem{Cui2008}
W.~Cui, H.~Zhou, H.~Qu, P.~C. Wong, and X.~Li, ``{G}eometry-based {E}dge
  {C}lustering for {G}raph {V}isualization,'' {\em {IEEE} Trans. Vis. Comput.
  Graph. ({TVCG})}, vol.~14, no.~6, 2008.

\bibitem{Holten2006}
D.~Holten, ``{H}ierarchical {E}dge {B}undles: {V}isualization of {A}djacency
  {R}elations in {H}ierarchical {D}ata,'' {\em {IEEE} Trans. Vis. Comput.
  Graph. ({TVCG})}, vol.~12, no.~5, 2006.

\bibitem{StolperPG14}
C.~D. Stolper, A.~Perer, and D.~Gotz, ``{P}rogressive {V}isual {A}nalytics:
  {U}ser-driven {V}isual {E}xploration of {I}n-progress {A}nalytics,'' {\em
  {IEEE} Trans. Vis. Comput. Graph. ({TVCG})}, vol.~20, no.~12, 2014.

\bibitem{KalininCZ14}
A.~Kalinin, U.~{\c{C}}etintemel, and S.~B. Zdonik, ``{I}nteractive {D}ata
  {E}xploration {U}sing {S}emantic {W}indows,'' in {\em {ACM} Conference on
  Management of Data ({SIGMOD})}, 2014.

\bibitem{Alagiannis2012}
I.~Alagiannis, R.~Borovica, M.~Branco, S.~Idreos, and A.~Ailamaki, ``{N}odb:
  {E}fficient {Q}uery {E}xecution on {R}aw {D}ata {F}iles,'' in {\em {ACM}
  Conference on Management of Data ({SIGMOD})}, 2012.

\bibitem{OlmaKAAA17}
M.~Olma, M.~Karpathiotakis, I.~Alagiannis, M.~Athanassoulis, and A.~Ailamaki,
  ``{S}lalom: {C}oasting through {R}aw {D}ata {V}ia {A}daptive {P}artitioning
  and {I}ndexing,'' {\em VLDB Endowment (PVLDB)}, vol.~10, no.~10, 2017.

\bibitem{ZoumpatianosIP14}
K.~Zoumpatianos, S.~Idreos, and T.~Palpanas, ``{I}ndexing for {I}nteractive
  {E}xploration of {B}ig {D}ata {S}eries,'' in {\em {ACM} Conference on
  Management of Data ({SIGMOD})}, 2014.

\bibitem{KarpathiotakisBAA14}
M.~Karpathiotakis, M.~Branco, I.~Alagiannis, and A.~Ailamaki, ``{A}daptive
  {Q}uery {P}rocessing on {R}aw {D}ata,'' {\em VLDB Endowment (PVLDB)}, vol.~7,
  no.~12, 2014.

\bibitem{TianALAMV17}
Y.~Tian, I.~Alagiannis, E.~Liarou, A.~Ailamaki, P.~Michiardi, and M.~Vukolic,
  ``{D}inodb: {A}n {I}nteractive-speed {Q}uery {E}ngine for {A}d-hoc {Q}ueries
  on {T}emporary {D}ata,'' {\em {IEEE} Trans. Big Data}, vol.~3, no.~3, 2017.

\bibitem{bcs15}
L.~Battle, R.~Chang, and M.~Stonebraker, ``{D}ynamic {P}refetching of {D}ata
  {T}iles for {I}nteractive {V}isualization,'' in {\em {ACM} Conference on
  Management of Data ({SIGMOD})}, 2016.

\bibitem{TauheedHSMA12}
F.~Tauheed, T.~Heinis, F.~Sch{\"{u}}rmann, H.~Markram, and A.~Ailamaki,
  ``{SCOUT:} {P}refetching for {L}atent {F}eature {F}ollowing {Q}ueries,'' {\em
  VLDB Endowment (PVLDB)}, vol.~5, no.~11, 2012.

\bibitem{JayachandranTKN14}
P.~Jayachandran, K.~Tunga, N.~Kamat, and A.~Nandi, ``{C}ombining {U}ser
  {I}nteraction, {S}peculative {Q}uery {E}xecution and {S}ampling in the {D}ice
  {S}ystem,'' {\em VLDB Endowment (PVLDB)}, vol.~7, no.~13, 2014.

\bibitem{ChanXGH08}
S.~Chan, L.~Xiao, J.~Gerth, and P.~Hanrahan, ``{M}aintaining {I}nteractivity
  {W}hile {E}xploring {M}assive {T}ime {S}eries,'' in {\em IEEE VAST}, 2008.

\bibitem{KhanSA14}
H.~A. Khan, M.~A. Sharaf, and A.~Albarrak, ``{D}ivide: {E}fficient
  {D}iversification for {I}nteractive {D}ata {E}xploration,'' in {\em Intl.
  Conf. on Scientific and Statistical Database Management ({SSDBM})}, 2014.

\bibitem{VartakHSMP16}
M.~Vartak, S.~Huang, T.~Siddiqui, S.~Madden, and A.~G. Parameswaran,
  ``{T}owards {V}isualization {R}ecommendation {S}ystems,'' {\em {SIGMOD}
  Record}, vol.~45, no.~4, 2016.

\bibitem{Key2012}
A.~Key, B.~Howe, D.~Perry, and C.~R. Aragon, ``{V}izdeck: {S}elf-organizing
  {D}ashboards for {V}isual {A}nalytics,'' in {\em {ACM} Conference on
  Management of Data ({SIGMOD})}, 2012.

\bibitem{EhsanSC16}
H.~Ehsan, M.~A. Sharaf, and P.~K. Chrysanthis, ``{M}uve: {E}fficient
  {M}ulti-objective {V}iew {R}ecommendation for {V}isual {D}ata
  {E}xploration,'' in {\em {IEEE} Intl. Conf. on Data Engineering ({ICDE})},
  2016.

\bibitem{MackinlayHS07}
J.~D. Mackinlay, P.~Hanrahan, and C.~Stolte, ``{S}how {M}e: {A}utomatic
  {P}resentation for {V}isual {A}nalysis,'' {\em {IEEE} Trans. Vis. Comput.
  Graph. ({TVCG})}, vol.~13, no.~6, 2007.

\bibitem{EURECOM+4380}
G.~A. {A}temezing and R.~{T}roncy, ``{T}owards a {L}inked-data {B}ased
  {V}isualization {W}izard,'' in {\em Workshop on Consuming Linked Data
  (COLD)}, 2014.

\bibitem{ThellmannGOS15}
K.~Thellmann, M.~Galkin, F.~Orlandi, and S.~Auer, ``{L}inkdaviz - {A}utomatic
  {B}inding of {L}inked {D}ata to {V}isualizations,'' in {\em Intl. Semantic
  Web Conf. ({ISWC})}, 2015.

\bibitem{VartakMPP14}
M.~Vartak, S.~Madden, A.~G. Parameswaran, and N.~Polyzotis, ``{SEEDB:}
  {A}utomatically {G}enerating {Q}uery {V}isualizations,'' {\em VLDB Endowment
  (PVLDB)}, vol.~7, no.~13, 2014.

\bibitem{WongsuphasawatM16}
K.~Wongsuphasawat, D.~Moritz, A.~Anand, J.~D. Mackinlay, B.~Howe, and J.~Heer,
  ``{V}oyager: {E}xploratory {A}nalysis {V}ia {F}aceted {B}rowsing of
  {V}isualization {R}ecommendations,'' {\em {IEEE} Trans. Vis. Comput. Graph.
  ({TVCG})}, vol.~22, no.~1, 2016.

\bibitem{WillsW10}
G.~Wills and L.~Wilkinson, ``{A}utovis: {A}utomatic {V}isualization,'' {\em
  Information Visualization}, vol.~9, no.~1, 2010.

\bibitem{MutluVT16}
B.~Mutlu, E.~E. Veas, and C.~Trattner, ``{V}izrec: {R}ecommending
  {P}ersonalized {V}isualizations,'' {\em ACM Transactions on Interactive
  Intelligent Systems (TiiS)}, vol.~6, no.~4, 2016.

\bibitem{GotzW09}
D.~Gotz and Z.~Wen, ``{B}ehavior-driven {V}isualization {R}ecommendation,'' in
  {\em Intl. Conference on Intelligent User Interfaces {(IUI)}}, 2009.

\bibitem{KandelPPHH12}
S.~Kandel, R.~Parikh, A.~Paepcke, J.~M. Hellerstein, and J.~Heer, ``{P}rofiler:
  {I}ntegrated {S}tatistical {A}nalysis and {V}isualization for {D}ata
  {Q}uality {A}ssessment,'' in {\em Intl. Working Conference on Advanced Visual
  Interfaces {(AVI)}}, 2012.

\bibitem{WM13}
E.~Wu and S.~Madden, ``{S}corpion: {E}xplaining {A}way {O}utliers in
  {A}ggregate {Q}ueries,'' {\em VLDB Endowment (PVLDB)}, vol.~6, no.~8, 2013.

\bibitem{Murray13}
S.~Murray, {\em {I}nteractive {D}ata {V}isualization for the {W}eb}.
\newblock O'Reilly Media, Inc., 2013.

\bibitem{book0024324}
M.~O. Ward, G.~G. Grinstein, and D.~A. Keim, {\em {I}nteractive {D}ata
  {V}isualization - {F}oundations, {T}echniques, and {A}pplications}.
\newblock A {K} Peters, 2010.

\bibitem{daglib0028506}
D.~A. Keim, J.~Kohlhammer, G.~P. Ellis, and F.~Mansmann, {\em {M}astering the
  {I}nformation {A}ge - {S}olving {P}roblems with {V}isual {A}nalytics}.
\newblock Eurographics Association, 2010.

\bibitem{Johnson2004}
C.~Johnson and C.~Hansen, {\em {V}isualization {H}andbook}.
\newblock Academic Press, Inc., 2004.

\bibitem{LandesbergerKSKWFF11}
T.~von Landesberger, A.~Kuijper, T.~Schreck, J.~Kohlhammer, J.~J. van Wijk,
  J.~Fekete, and D.~W. Fellner, ``{V}isual {A}nalysis of {L}arge {G}raphs:
  {S}tate-of-the-art and {F}uture {R}esearch {C}hallenges,'' {\em Comput.
  Graph. Forum ({CGF})}, vol.~30, no.~6, 2011.

\bibitem{HermanMM00}
I.~Herman, G.~Melan{\c{c}}on, and M.~S. Marshall, ``{G}raph {V}isualization and
  {N}avigation in {I}nformation {V}isualization: {A} {S}urvey,'' {\em {IEEE}
  Trans. Vis. Comput. Graph. ({TVCG})}, vol.~6, no.~1, 2000.

\bibitem{BattistaETT99}
G.~{Di Battista}, P.~Eades, R.~Tamassia, and I.~G. Tollis, {\em {G}raph
  {D}rawing: {A}lgorithms for the {V}isualization of {G}raphs}.
\newblock Prentice-Hall, 1999.

\bibitem{DadzieP17}
A.~Dadzie and E.~Pietriga, ``{V}isualisation of {L}inked {D}ata - {R}eprise,''
  {\em Semantic Web}, vol.~8, no.~1, 2017.

\bibitem{MarieG14a}
N.~Marie and F.~L. Gandon, ``{S}urvey of {L}inked {D}ata {B}ased {E}xploration
  {S}ystems,'' in {\em Intl. Workshop on Intelligent Exploration of Semantic
  Data}, 2014.

\bibitem{DR11}
A.~Dadzie and M.~Rowe, ``{A}pproaches to {V}isualising {L}inked {D}ata: {A}
  {S}urvey,'' {\em Semantic Web}, vol.~2, no.~2, 2011.

\bibitem{LiuMWBP17}
S.~Liu, D.~Maljovec, B.~Wang, P.~Bremer, and V.~Pascucci, ``{V}isualizing
  {H}igh-dimensional {D}ata: {A}dvances in the {P}ast {D}ecade,'' {\em {IEEE}
  Trans. Vis. Comput. Graph. ({TVCG})}, vol.~23, no.~3, 2017.

\bibitem{WohlfartABM08}
E.~Wohlfart, W.~Aigner, A.~Bertone, and S.~Miksch, ``{C}omparing {I}nformation
  {V}isualization {T}ools {F}ocusing on the {T}emporal {D}imensions,'' in {\em
  Intl. Conference on Information Visualisation {(IV)}}, 2008.

\end{thebibliography}

\documentclass{svmult}

\usepackage[makeroom]{cancel}
\usepackage{pifont}
\usepackage{marvosym}

 \usepackage[justification=centering]{caption} 
\usepackage[table]{xcolor}
 \usepackage[caption=false]{subfig}
\usepackage{color}
\usepackage{multirow}
 \usepackage{graphicx}
\usepackage[linesnumbered, vlined, ruled,algo2e]{algorithm2e} 
\usepackage{booktabs}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{mathtools}

\usepackage{amsthm}
\usepackage{soul}
\usepackage{etex}
 \usepackage{relsize}
\usepackage{epstopdf}
\usepackage{multirow}
\usepackage{url}
\usepackage{tablefootnote}
\usepackage{setspace}
 \usepackage{hyperref}  
\hypersetup{ 
    colorlinks,%
    citecolor=black,%
    filecolor=black,%
    linkcolor=black,%
    urlcolor=black,%
}
\usepackage[nocompress]{cite}

\newcommand{\eat}[1]{}

\definecolor{dark-gray}{gray}{0.2}

\newcounter{corollcnt}

\newcounter{claimcnt}

\newcounter{remarkcnt}

\newcounter{proposcnt}

\newcounter{lemmcnt}
{\medskip\par}

\newcounter{ex}

\newcommand{\myFontP}[1]{{\fontfamily{ppl}\selectfont #1}} 

\newcolumntype{'}{!{\vrule width 0.6pt}}

\SetKw{Logand}{and}	
\SetKw{Logor}{or}
\SetKw{Lognot}{not}	

\SetKw{mybreak}{break}	
\SetKw{mycontinue}{continue}	

\SetKwInput{KwVar}{Variables}
\SetKwInput{KwPar}{Parameters}

\SetKwData{myEOF}{EOF}	
\SetKwData{myfalse}{false}
\SetKwData{mytrue}{true}
\SetKwData{mynull}{null}
\SetKwData{myvalid}{\texttt{valid}}
\SetKwData{myfeasible}{\texttt{feasible}}
\SetKwData{myupdated}{\texttt{updated}}
\SetKwData{mynotupdated}{\scalebox{0.9}[1]{ \texttt{prtl  \!\! \!\!\!updated}}}

\SetKwData{myempty}{empty}
\SetKwBlock{myblock}{beginnikos}{end} 
\SetAlgoInsideSkip{} 
\SetAlgoSkip{} 
\SetVlineSkip {0.5mm} 
\SetAlgoCaptionSeparator{.} 

\newcommand{\mycomment} [1] 	{\tiny{\textcolor{dark-gray}{\myFontP{{#1}}}}} 
\SetKwComment{Comment}{\mycomment{/\!\!/}}{}
 \DontPrintSemicolon
\SetAlgoCaptionLayout{small}
\SetProcNameSty{small}
\SetProcArgSty{small}

\graphicspath{{figures/}{plots/}}

 \hypersetup{ 
    colorlinks,%
    citecolor=black,%
    filecolor=black,%
    linkcolor=black,%
    urlcolor=black,%
pdftitle={ Big Data Visualization Tools Systems A Survey of the State of the Art and Challenges},
pdfauthor={Nikos Bikakis},
pdfsubject={Big Data Visualization Tools  Systems A Survey of the State of the Art and Challenges},
pdfproducer={pdfbikTeX-1.1},
pdfcreator={bikTeX},
pdfkeywords={encyclopedia, survey, visualization systems, big data, visualization tools, exploration, state of the art, 
Visual exploration, big data challenges, Interactive visualization, scalability, Information visualization, Visual analytics, Exploratory data analysis, big data era, challenges, in situ querying,  study, data exploration, in situ exploration,  state-of-the-art, large databases,  web of data, study,  large data visualization, visualizing very large datasets,  incremental visual exploration, progressive visualization, incremental visualization, multiresolution visualization, multiscale visualization, multilevel visualization, hierarchical visualization, adaptive visualization, data summarization, data abstraction, binning, binning aggregation, visualization recommendations, Steerable Exploration, preferences,  visual aggregation, scalable visual analytics,  HETree, Faceted search}
}

\begin{document}
%


\title*{Big Data Visualization Tools \thanks{\textbf{This article appears in \textit{Encyclopedia of Big Data Technologies}, Springer, 2018}}}
\author{Nikos Bikakis \\ ATHENA Research Center, Greece 
}
 \authorrunning{Nikos Bikakis} 


\maketitle

\section{Synonyms}
Visual   exploration;
Interactive  visualization;
Information visualization;
Visual analytics;
Exploratory data analysis.

 \section{Definition}
 

\textit{Data visualization} is the presentation of data in a pictorial or graphical format,  
and a \textit{data visualization tool} is the software that generates this presentation.
Data visualization provides users with intuitive means to interactively explore and analyze data, enabling them to 
effectively identify interesting patterns, infer correlations and causalities, and supports sense-making activities.

\section{Overview}
\label{sec:over}


Exploring, visualizing and analysing data is a core
task for data scientists and analysts in numerous applications.
\textit{Data visualization}\footnote{Throughout the article,   terms \textit{visualization} and \textit{visual exploration}, as well as terms \textit{tool} and \textit{system} are used interchangeably.}\cite{idvBook_v2} provides intuitive
ways for the users to interactively explore and analyze data, 
enabling them to effectively identify interesting patterns, infer correlations
and causalities, and support sense-making activities.

 The Big Data era has realized the availability of a great amount of \textit{massive} datasets that are \textit{dynamic}, \textit{noisy} and   \textit{heterogeneous} in nature.
The level of difficulty  in transforming a data-curious user 
into someone who can access and analyze that data is 
even more burdensome now for a great number of users 
with little or no support and expertise on the data processing part.
 Data visualization has become a major  research challenge 
involving several issues related to data storage, querying, indexing,  
visual presentation, interaction,  personalization 
   \cite{bs16,Shneiderman08,HeerK12b, MortonBGM14,WuBM14,GGL15,IdreosPC15,HeerS12}.   

Given the above,
 modern visualization and exploration systems should  effectively and efficiently handle the following aspects.
 
 \begin{itemize}
\item 
 \textit{Real-time Interaction}. 
Efficient and scalable  techniques  should support the interaction with 
billion objects  datasets,   while maintaining  the system response in 
the range of    a few milliseconds.

\item 
 \textit{On-the-fly Processing}. 
Support of on-the-fly  visualizations over large and dynamic sets of volatile raw (i.e., not preprocessed) data is required. 

\item 
 \textit{Visual Scalability}.  
 Provision of effective data abstraction mechanisms is necessary for addressing problems related to visual information overloading (a.k.a.\ overplotting). 

\item 
\textit{User Assistance and Personalization}. 
Encouraging user comprehension and offering customization capabilities to different user-defined exploration scenarios and preferences according to the analysis needs   are important features.

 \end{itemize}

%
%
%

\section{Visualization in Big Data Era}
\label{sec:bigvis}

   This section discusses the basic concepts related to Big Data visualization.
   First, the   limitations of traditional visualization systems are outlined.  
 Then, the basic characteristics of data visualization  in the context of   Big Data era  are presented.
	Finally, the  major prerequisites and challenges that should be addressed by   modern exploration
and visualization systems are discussed.

 \subsection{Traditional Systems}
Most \textit{traditional  exploration and visualization systems} 
 cannot handle the size of many contemporary datasets. 
They restrict themselves to dealing with \textit{small} dataset sizes, which can be easily handled and analysed with conventional data management and visual explorations techniques. 
Further, they operate in an \textit{offline} way, limited to accessing
\textit{preprocessed} sets of \textit{static} data.

%


\subsection{Current Setting}
On the other hand, nowadays, the \textit{Big Data era} has 
made available large numbers of \textit{very big} datasets, that are often \textit{dynamic} and characterized by high \textit{variety} and \textit{volatility}.
For example, in several cases 
 (e.g., scientific databases), new data   constantly arrive (e.g., on a   daily/hourly  basis);
 in other cases, data sources offer query or API endpoints  for online access and updating.
Further,  nowadays,  an increasingly  large number of \textit{diverse users} (i.e., users with different preferences or skills)  explore and analyze data  in a plethora of  \textit{different scenarios}.

\subsection{Modern Systems}
\textit{Modern  systems} should be able to 
efficiently handle \textit{big dynamic} datasets, 
operating on machines with limited computational and memory resources (e.g.,  laptops).
The dynamic nature of nowadays data (e.g., stream data),  
 hinders the application of a preprocessing phase, such as  traditional database loading and indexing.
Hence, systems should provide \textit{on-the-fly processing} over 
large sets of raw data.

Further, in conjunction with performance issues, modern systems have to 
address challenges related to \textit{visual presentation}. 
Visualizing a large number of data objects is a challenging task; 
modern systems have to  ``\textit{squeeze a billion records into a million pixels}" \cite{Shneiderman08}. 
Even in  small datasets,  offering a dataset \textit{overview}  may be extremely difficult; 
in both cases,  \textit{information overloading} (a.k.a.\ overplotting)  is a common issue.
Consequently, a basic requirement of   modern systems  
is to effectively support  \textit{data abstraction} 
over   enormous numbers of data objects.  

Apart from the aforementioned requirements, 
modern systems  must 
also satisfy
the diversity of \textit{preferences} and \textit{requirements} posed by different \textit{users}  and \textit{tasks}. 
Modern systems should provide the user with the ability to customize the exploration experience based on her preferences and the individual  requirements 
of each  examined task. 
Additionally, systems should
automatically adjust their parameters by taking into account the
\textit{environment setting} and \textit{available resources}; e.g., screen resolution/size, available memory.



%

%
%


\section{Systems  and  Techniques}
\label{sec:methods}

This section presents how state-of-the-art approaches from   
Data Management and Mining, Information Visualization and Human-Computer Interaction communities attempt to handle 
the challenges that arise in the Big Data era.

\subsection{Data Reduction}
In order to handle and visualize large datasets,  modern systems have to 
deal with information overloading issues. 
Offering \textit{visual scalability} are  crucial  in Big Data visualization. 
Systems should  provide efficient and effective abstraction and summarisation mechanisms.
In this direction, a large number of   systems 
use \textit{approximation techniques} (a.k.a.\ \textit{data reduction} techniques),
in which  abstract sets of data are computed. 
Considering the existing approaches, most of them are  based on: 
(1)~\textit{sampling} and \textit{filtering}
\cite{FisherPDs12,ParkCM15,AgarwalMPMMS13,ImVM13,BattleSC13}
and/or 
(2)~\textit{aggregation} (e.g., binning, clustering) \cite{EF10, bsps15,JugelJM15,LiuJH13,LinsKS13,GodfreyGLR16}.

 \subsection{Hierarchical Exploration}

Approximation techniques are often defined in a hierarchical manner \cite{EF10, bsps15,GodfreyGLR16,LinsKS13}, allowing users to explore data in different levels of detail (e.g., hierarchical aggregation).

\textit{Hierarchical approaches}\footnote{sometimes also referred as \textit{multilevel}}
 allow the visual exploration of
very large datasets in multiple levels, offering both an
overview, as well as an intuitive and effective
way for finding specific parts within a dataset. Particularly,
in hierarchical approaches, the user first obtains
an \textit{overview} of the dataset before proceeding to data exploration
operations (e.g., \mbox{roll-up}, \mbox{drill-down}, zoom, filtering) and finally retrieving \textit{details}
about the data. 
Therefore, hierarchical approaches directly
support the visual information seeking mantra
``\textit{overview first, zoom and filter, then details on demand}'' \cite{Shneiderman96}.
Hierarchical approaches can also effectively address
the problem of information overloading as they adopt 
approximation techniques.

Hierarchical techniques have been extensively used in 
large \textit{graphs visualization},  where the graph is recursively decomposed into 
smaller sub-graphs that form a hierarchy of abstraction layers.
In most cases, the hierarchy is constructed by exploiting \textit{clustering} and \textit{partitioning} methods
 \cite{RodriguesTPTTF13,AbelloHK06,Auber04,TominskiAS09}.
In other works, the hierarchy is defined with \textit{hub-based} \cite{LinCTWKC13} and \textit{density-based} \cite{ZinsmaierBDS12} techniques. \cite{ArchambaultMA08} supports \textit{ad-hoc hierarchies} which are manually defined by the users. 
Differents approaches have been adopted in \cite{SundaraAKDWCS10,BikakisLKG16},
 where  sampling techniques have been exploited. 
Other works adopt edge bundling techniques
 which  aggregate graph edges to bundles  \cite{Gansner2011,Ersoy2011,Phan2005,Lambert2010,Cui2008,Holten2006}.

%

\subsection{Progressive {Results}}

%
Data exploration requires real-time system's response.
However, 
computing \textit{complete results} over large (unprocessed) 
datasets may be extremely costly and in several cases \textit{unnecessary}. 
Modern systems should \textit{progressively} return 
\textit{partial} and preferably \textit{representative} results, as soon as possible.

\textit{Progressiveness} can significantly improve efficiency in \textit{exploration scenarios}, 
where it is common that 
users  attempt to find something interesting without  
knowing what exactly they are searching for
beforehand.
In this case, users perform a sequence of operations (e.g., queries), where the result of each operation determines the formulation of the  next operation. 
%
%
{In systems where progressiveness is supported, 
in each operation, 
after inspecting the already produced results, 
the user is able to interrupt the execution and define the next  operation, 
without waiting the exact result to be computed.
}

In this context, 
several systems adopt  \textit{progressive techniques}.
In these techniques the results/visual elements are computed/constructed incrementally based on  user interaction or as time progresses  \cite{StolperPG14, bsps15,KalininCZ14}.
Further, numerous recent systems integrate incremental and approximate techniques.
In these cases, approximate results are computed incrementally over progressively larger samples of the data \cite{FisherPDs12,AgarwalMPMMS13,ImVM13}.

 

 \subsection{Incremental and Adaptive Processing}
 The dynamic setting established nowadays 
hinders (efficient) data preprocessing in  modern systems.
 Additionally, it is common in exploration scenarios that 
only a small fragment of the input data to be  accessed by the user.

\textit{In situ data exploration} \cite{Alagiannis2012,OlmaKAAA17,bsps15,ZoumpatianosIP14,KarpathiotakisBAA14,TianALAMV17}  %
is a recent trend, which aims at enabling on-the-fly  exploration over large and dynamic sets of  data, 
without (pre)processing (e.g., loading, indexing) the whole dataset.  
In these systems, \textit{incremental} and \textit{adaptive} processing and indexing techniques are used, in which small parts of data are processed  incrementally ``following'' users' interactions.

\subsection{Caching and Prefetching}
Recall that, in exploration scenarios, a sequence of operations is performed and, in most cases, each operation is driven by the previous one.
In this setting, \textit{caching} and/or \textit{prefetching} the sets of data that are likely to be accessed by the user in the near future can significantly reduce the response time 
\cite{bcs15,TauheedHSMA12,KalininCZ14,JayachandranTKN14,bsps15,ChanXGH08,KhanSA14}.  %
Most of these approaches use prediction techniques  which exploit several factors 
(e.g., user behavior, user profile, use case) in order to determine the upcoming user interactions.

\subsection{{User Assistance}}
The large amount of available information makes
it difficult for users to manually explore and analyze data.
Modern systems should provide mechanisms that assist 
the user and reduce the effort needed on their part, considering 
the diversity of {preferences} and {requirements} posed by different {users}  and {tasks}.

Recently,  several approaches have been developed in the context of \textit{visualization recommendation} \cite{VartakHSMP16}.
These approaches recommend the most suitable  visualizations in order to assist users 
throughout the analysis process. 
Usually, the recommendations take into account several factors, such as
data characteristics, examined task, user preferences and behavior, etc.

Especially considering data characteristics, there are several systems that recommend the most suitable visualization technique (and parameters) based on the type, attributes, distribution, or cardinality of the input data
\cite{Key2012,EhsanSC16,MackinlayHS07,bsps15,EURECOM+4380,ThellmannGOS15}.
%
 %
In a similar context, some systems assist users by recommending certain visualizations 
that  reveal  surprising and/or interesting data \cite{VartakMPP14,WongsuphasawatM16,WillsW10}. 
Other approaches provide  visualization recommendations based on user behavior and preferences \cite{MutluVT16,GotzW09}.
%
Finally, systems provide recommendations and explanations 
regarding data trends and anomalies \cite{KandelPPHH12,WM13}.

%
%

\section{Examples of Applications}
\label{sec:appl}
    
 Visualization techniques are of great importance in a wide range of application areas in the Big Data era.  The volume, velocity, heterogeneity and complexity of available data make it extremely  difficult for humans to explore and analyze data.
Data visualization enables users to perform a series of analysis tasks that are not always possible with common data analysis techniques \cite{daglib0028506}.

Major application domains for data visualization and analytics are \textit{Physics} and \textit{Astronomy}.
Satellites and telescopes  collect daily massive and dynamic streams of data.  
Using traditional analysis techniques, astronomers are able to identify noise, patterns and similarities. 
On the other hand, visual analytics can enable astronomers to identify unexpected phenomena and perform several complex operations, which are not are feasible by traditional analysis approaches.

Another application domain is \textit{atmospheric sciences} like \textit{Meteorology}
and \textit{Climatology}. In this domain high volumes of data are collected from sensors and satellites on a daily basis. Storing these data over the years results in massive amounts of data that have to be analyzed. Visual analytics can assist scientists to perform core tasks, such as climate factors correlation analysis, event prediction, etc. 
Further, in this domain, visualization systems are used in several scenarios in order capture real-time phenomena, such as, hurricanes, fires, floods, and tsunamis.

In the domain of \textit{Bioinformatics}, visualization techniques are exploited in numerous tasks. For example, analyzing the large amounts of biological data   produced by DNA sequencers is extremely challenging. Visual techniques can help biologist to gain insight and identify interesting ``areas’’ of genes on which to performs their experiments. 

In the Big Data era, visualization techniques are extensively used in the \textit{business intelligence} domain. \textit{Finance markets} is one application area, where visual analytics allow to monitor markets, identify trends  and perform predictions. 
Besides, \textit{market research} is also an application area. Marketing agencies and in-house marketing departments analyze a plethora of diverse sources (e.g., finance data, customer behavior, social media). Visual techniques are exploited to realize task such as, identifying trends, finding emerging market opportunities, finding influential users and communities, optimizing operations (e.g., troubleshooting of products and services), business analysis and development (e.g., churn rate prediction, marketing optimization).

\section{Further Reading}
\label{sec:reading}

 The literature on visualization is extensive, covering a large range of fields and many decades. 
Data visualization is discussed in a great number of recent introductory-level textbooks, such as \cite{idvBook_v2,Murray13,book0024324,daglib0028506,Johnson2004}.

Also, there are various articles discussing Big Data visualization; see   \cite{Shneiderman08,HeerK12b, MortonBGM14,WuBM14,HeerS12}.   
Surveys of Big Data visualization systems can be found at  \cite{GGL15,bs16,IdreosPC15}.

In what follows we provide some surveys/studies related to issues discussed in this article: 
 \begin{itemize}
\item 
\textit{Graph visualization} \cite{LandesbergerKSKWFF11,HermanMM00,BattistaETT99}
\item 
\textit{Hierarchical exploration} \cite{EF10}
\item 
\textit{Visualization recommendations} \cite{VartakHSMP16}
\item 
\textit{Linked and Web data visualization}  \cite{bs16,DadzieP17,MarieG14a,DR11}
\item 
\textit{High-dimensional data visualization} \cite{LiuMWBP17}
\item 
\textit{Temporal data visualization} \cite{WohlfartABM08}
\end{itemize}

%

Some of the major workshops and symposiums focusing on Big Data visualization include:
 \begin{itemize}
\item
\textit{Workshop on Big Data Visual Exploration and Analytics} (BigVis)
\item
\textit{Symposium on Big Data Visual Analytics} (BDVA)
\item
\textit{Big Data Analysis and Visualization} (LDAV)
\item
\textit{Workshop on Data Mining Meets Visual Analytics at Big Data era} (DAVA)
\item
\textit{Workshop on Human-In-the-Loop Data Analytics} (HILDA)
\item 
\textit{Workshop on Data Systems for Interactive Analysis} (DSIA)
\item
\textit{Workshop on Immersive Analytics:  Exploring Future Interaction and Visualization Technologies for Data Analytics}
\end{itemize}

%

Finally, there is a great deal of information regarding visualization tools available in the Web. 
We mention {\myFontP{dataviz.tools}}\footnote{\href{http://dataviz.tools}{{http://dataviz.tools}}}  
and  \myFontP{datavizcatalogue}\footnote{\href{http://www.datavizcatalogue.com}{{www.datavizcatalogue.com}}} which are catalogs containing a large number of visualization tools, libraries and resources.

\section{Cross-References}
  \begin{itemize}
\item Visualization
\item Visualization Techniques
\item Visualizing Semantic Data 
\item Graph exploration and search
  \end{itemize}

 %


\bibliographystyle{ieeetr}


\bibliography{biblio}

\end{document}